A copper calorimeter of mass 120g contains 70g of ice at 0degree c. What mass of steam at 100degree c must be passed into the calorimeter to raise a temperature to 40 degree c.
1
Expert's answer
2020-05-25T11:05:12-0400

Heat gained by the copper:


"q_1=m_cC_cΔT

 = (120) (0.40 ) (40 − 0)

= 1920\ J"

Heat gained by the water:


"q_2=m_wC_wΔT

 = (70) (4.2 ) (40 − 0)

= 11760\ J"

Heat gained by the ice:


"q_3 = m_iL + m_iC_wΔT

\\ = (10) (320 ) + (10 ) (4.2 ) (40-0 )

 = 4880\ J"

Heat lost by the steam:


"q _4= mL_s + mC_wΔT

 \\= m (2200) + m (4.2 ) (100 − 40)

 = 2452 m"

"1920 + 11760 + 4880 = 2452 m"

"m = 7.6\ g"